# A Transferable Physics-Informed Framework for Battery Degradation Diagnosis, Knee-Onset Detection and Knee Prediction

### v1

@@ Abstract
The techno-economic and safety concerns of battery capacity knee occurrence call for developing online knee detection and prediction methods as an advanced battery management system (BMS) function. To address this, a transferable physics-informed framework that consists of a histogram-based feature engineering method, a hybrid physics-informed model, and a fine-tuning strategy, is proposed for online battery degradation diagnosis and knee-onset detection. The hybrid model is first developed and evaluated using a scenario-aware pipeline in protocol cycling scenarios and then fine-tuned to create a local model deployed in a dynamic cycling scenario. A 2D histogram-based feature set is found to be the best choice in both source and target scenarios. The fine-tuning strategy is proven to be effective in improving battery degradation mode estimation and degradation phase detection performance in the target scenario. Again, a strong linear correlation was found between the identified knee-onset and knee points. As a result, advanced BMS functions, such as online degradation diagnosis and prognosis, online knee-onset detection and knee prediction, aging-aware battery classification, and second-life repurposing, can be enabled through a battery performance digital twin in the cloud.
@@
A 2D histogram-based method to enable online battery diagnosis and knee-onset detection.
@@
A fine-tuning strategy to create a local model deployed in the dynamic cycling scenario.
@@ 1 Introduction
Contribution: The goal of this work is to fill the gaps indicated above by proposing a transferable physics-informed framework for online battery degradation diagnosis, knee-onset detection, and knee prediction. Specifically, the model takes histograms aggregated from time-series voltage and current data, which are easy to acquire in the onboard BMS. In addition, its performance can be generalized to unseen battery usage scenarios at a satisfactory level using a small amount of labeled data in a target scenario. Our key results and contributions are as follows:
@@
    To start with, batteries are classified according to which degradation phase they are in and then battery next-stage decision-making can be enabled in their first lives. Specifically, Phase 1 is defined as the period from the beginning of life to the knee-onset point in which batteries may continue their usage in first-life applications; Phase 2 is defined as the period from the knee-onset point to the knee point in which batteries may either continue their usage in first-life applications or be repurposed to second-life applications in which the knee occurrence may be avoided; Phase 3 is defined as the period beyond the knee point in which batteries may potentially be repurposed to very mild second-life applications in which the knee occurrence can be stopped or be recycled.
@@
    A transferable physics-informed framework is proposed, which consists of a histogram-based feature engineering method, a hybrid physics-informed model, and a fine-tuning strategy. In total, 5 histogram-based feature sets are generated for online battery degradation mode estimation. The hybrid physics-informed model consists of a deep hidden physics model (DeepHPM) for battery degradation mode estimation and an XGBoost model for battery degradation phase detection. This hybrid model is first developed and evaluated using a scenario-aware pipeline in the protocol cycling scenarios and then fine-tuned to create a local model deployed in a dynamic cycling scenario.
@@
    The hybrid model using 2D histogram-based features is found to be the best choice in both source and target scenarios. The fine-tuning strategy was proven to be effective in improving battery degradation mode estimation and phase detection performance in the target scenario. With degradation phases detected with high accuracy, online prediction of battery capacity knee points can also be achieved by leveraging the strong linear correlations identified between knee-onset and knee points.
@@ 2.1 Related definitions
A degradation mode is a degradation mechanism that has a unique and measurable effect on, for example, the capacity, impedance, and open circuit voltage (OCV) of lithium-ion cells and electrodes, caused by one or multiple interacting degradation mechanisms. Three degradation modes are considered here, loss of lithium inventory (LLI), loss of active material at the negative electrode (LAM_NE), and loss of active material at the positive electrode (LAM_PE). Notably, another degradation mode, i.e., conductivity loss (CL), describes ohmic resistance increase.
Definition 2.3 (Knee-onset, knee, and degradation phases [37]).
@@ 2.2 Aging-aware battery classification
Lithium-ion batteries are designed, manufactured, and used for a wide range of applications, such as portable electronics, electric vehicles (EVs), and stationary energy storage systems. As a result of complex interactions between multiple degradation mechanisms that are influenced by operating conditions (e.g., charge/discharge current, temperature, state-of-charge (SoC) window, etc.), batteries can exhibit capacity fade curves that are linear, sublinear, superlinear or a combination of three [21]. In particular, superlinear capacity fade infers a knee on the capacity fade curve, which may significantly shorten a battery’s lifetime and pose safety risks [38]. Therefore, batteries with knee occurrence are recommended to retire immediately for safety concerns [39]. As a result, retired EV batteries with knee occurrence will have no second-life value. However, if we can classify EV batteries during their first-life applications, and then repurpose them to well-controlled second-life applications in which the knee occurrence can be delayed or even avoided, then enormous quantities of EV batteries could potentially have substantial second-life value with guaranteed safety.
@@ 2.3 Histogram-based feature engineering
This histogram-based feature engineering method consists of two steps, i.e., first determining the variable bounds; and then extracting features from time-series usage data. We assume that voltage and current measurements are available for each cell in a battery pack, we will therefore focus on features extracted from the time series of these two variables in this work. Specifically, we first determine the variable bounds, for which a histogram and a cumulative histogram are generated from the measured voltage and current data. An example of the histogram and the cumulative one generated from the voltage data can be seen in Fig. 2a and 2b, respectively. Similarly, an example of the histogram and the cumulative one generated using the current data can be seen in Fig. 3a and 3b, respectively. Each bar in the histograms represents the time spent within a specific voltage or current range. The 1st, 33rd, 67th, and 99th percentiles of each variable are given in Table 1. Based on these bounds, a voltage-based 3-feature set was selected as the optimal set for online capacity estimation by Greenbank and Howey [15]. Although battery capacity fade is attributed to the growth of internal degradation modes, this feature set may not be the best feature set for online battery degradation mode estimation. Moreover, time spent in extreme ranges, for example, ranges outside charge/discharge cut-off voltages, should also be monitored due to its evident degradation effects [42]. As listed in Table 2, 5 feature sets are generated for online battery degradation mode estimation.
@@
Feature set Input feature Voltage-based 3-feature set Time spent between 1st and 33rd voltage percentiles (i.e., $V_{12}$) Time spent between 33rd and 67th voltage percentiles (i.e., $V_{23}$) Calendar time (i.e., $t$) Voltage-based 5-feature set Time spent less than 1st voltage percentile (i.e., $V_{01}$) Time spent between 1st and 33rd voltage percentiles (i.e., $V_{12}$) Time spent between 33rd and 67th voltage percentiles (i.e., $V_{23}$) Time spent greater than 67th voltage percentile (i.e., $V_{34}$) Calendar time (i.e., $t$) Current-based 3-feature set Time spent between 1st and 33rd current percentiles (i.e., $I_{12}$) Time spent between 33rd and 67th current percentiles (i.e., $I_{23}$) Calendar time (i.e., $t$) Current-based 5-feature set Time spent less than 1st current percentile (i.e., $I_{01}$) Time spent between 1st and 33rd current percentiles (i.e., $I_{12}$) Time spent between 33rd and 67th current percentiles (i.e., $I_{23}$) Time spent greater than 67th current percentile (i.e., $I_{34}$) Calendar time (i.e., $t$) Current-Voltage 17-feature set Time spent less than 1st current percentile and in 4 voltage ranges (i.e., $I_{01}V_{01}$, $I_{01}V_{12}$. $I_{01}V_{23}$. $I_{01}V_{34}$) Time spent between 1st and 33rd current percentiles and in 4 voltage ranges (i.e., $I_{12}V_{01}$, $I_{12}V_{12}$, $I_{12}V_{23}$, $I_{12}V_{34}$) Time spent between 33rd and 67th current percentiles and in 4 voltage ranges (i.e., $I_{23}V_{01}$, $I_{23}V_{12}$, $I_{23}V_{23}$, $I_{23}V_{34}$) Time spent greater than 67th current percentile and in 4 voltage ranges (i.e., $I_{34}V_{01}$, $I_{34}V_{12}$, $I_{34}V_{23}$, $I_{34}V_{34}$) Calendar time (i.e., $t$)
@@ 2.4 Hybrid model architecture design
$u=f(t,\bm{x}),$ (1)
where $t\in\mathbb{R}$ denotes time, $\bm{x}\in\mathbb{R}^{m}$ denotes an input vector, and $u\in\mathbb{R}$ denotes one of the three degradation modes. The degradation mode growth rate can be described using a nonlinear partial differential equation (PDE) in the general form
@@
where $g(\cdot)$ is a nonlinear function of time $t$, the input vector $\bm{x}$, solution $u$, and its derivatives with respect to the input vector, for example, $u_{\bm{x}}=[\partialderivative{u}{x_{1}},\partialderivative{u}{x_{2}},\ldots,\partialderivative{u}{x_{m}}]^{T}$. The function $g(\cdot)$ comprises battery internal degradation dynamics and can represent different forms of degradation, such as linear, sublinear, superlinear, or combination of the three [21]. However, the explicit form of $g(\cdot)$ is unknown and almost impossible to obtain. Inspired by the deep hidden physics model (DeepHPM) proposed by Raissi et al. [43], we approximate the function $f(\cdot)$ in Eqn. (1) and the nonlinear function $g(\cdot)$ in Eqn. (2) with two neural networks (NNs) and define a DeepHPM $\mathcal{H}$ to model battery degradation mode:
@@
where $\mathcal{F}(\cdot)$ denotes the surrogate NN that approximates the hidden solution of the dynamical models, $\mathcal{G}(\cdot)$ denotes the dynamic NN that approximates the battery degradation dynamics, and $\partialderivative{\mathcal{F}(t,\bm{x};\Phi)}{t}$ denotes the partial derivatives of the surrogate NN $\mathcal{F}(\cdot)$ with respect to $t$. The input vector $\bm{x}$ can be one of 5 feature sets, excluding the calendar time $t$ in Table 2. Notably, we only consider the first-order partial derivatives in the dynamic NN, to achieve a good trade-off between accuracy and complexity. The parameters $\Phi$ of the surrogate NN $\mathcal{F}(\cdot)$ and $\Theta$ of the dynamic NN $\mathcal{G}(\cdot)$ can be learned by minimizing the sum of squared errors by the loss function
@@
where $n$ is the number of training samples and $\mathcal{H}_{t}=\partialderivative{\mathcal{H}}{t}$. The data loss term $\mathcal{L}_{u}$ aims to find the parameters of the surrogate NN $\mathcal{F}(\cdot)$ so that it fits the data, while the PDE loss term $\mathcal{L}_{\mathcal{H}}$ aims to find the parameters of the dynamic NN $\mathcal{G}(\cdot)$ so that it satisfies the PDE defined by Eqn. (2) at the evaluated points $(t_{i},\bm{x}_{i})$. Moreover, it has been empirically demonstrated that embedding the gradient of the PDE residual into the loss function could further improve the accuracy of the DeepHPM [44]. Therefore the PDE gradient loss term $\mathcal{L}_{\mathcal{H}_{t}}$ that aims to reduce fluctuations and makes the PDE residual closer to zero is also added here. During the training process, the derivatives of the surrogate NN w.r.t. time $t$ and input $\bm{x}$, and the derivatives of the DeepHPM function w.r.t. time $t$ are evaluated using automatic differentiation [45].
With the availability of battery degradation modes estimated by the DeepHPM, the degradation phase of a battery can be detected. Mathematically, the degradation phase detection can be formulated as a multi-class classification problem to classify each battery into one of the 3 possible classes (see Subsection 2.2) given its estimated degradation modes and calendar time. As an efficient and scalable machine learning system for tree boosting, proposed by Chen and Guestrin [46], XGBoost has demonstrated outstanding performance in battery state of health estimation [47] [48], and remaining useful life prediction [49]. Thus, it is reasonable to explore its potential for addressing the classification problem here.
@@
$\mathcal{L}^{(t)}=\sum^{n}_{i=1}l(y_{i},\hat{y}^{(t-1)}_{i}+h_{t}(\bm{x}_{i}))+\omega(h_{t})+\mathrm{constant},$ (8)
where $n$ is the number of training samples, $y_{i}$ is the value of the $i$-th sample, $\hat{y}^{t-1}_{i}$ is the prediction of the $i$-th sample up to the $(t-1)$-th iteration, $h_{t}(\bm{x}_{i})$ is the output of the $t$-th tree for the $i$-th sample. In our case, the XGBoost model is to approximate a function that predicts the class $y\in\{1,2,3\}$ given an input vector $\bm{x}=[u_{1},u_{2},u_{3},t]^{T}\in\mathbb{R}^{4}$. Thus, the general loss function in Eqn. (8) must be modified to account for multi-class classification as
$\mathcal{L}^{(t)}=\sum^{n}_{i=1}\sum^{3}_{k=1}l(y^{k}_{i},\hat{y}^{k,(t-1)}_{i}+h^{k}_{t}(\bm{x}_{i}))+\omega(h_{t})+\mathrm{constant},$ (9)
@@
The resulting hybrid physics-informed model is illustrated in Fig. 4.
Figure 4: Hybrid physics-informed model.

### v2

@@ Abstract
The techno-economic and safety concerns of battery capacity knee occurrence call for developing online knee detection and prediction methods as an advanced battery management system (BMS) function. To address this, a transferable physics-informed framework that consists of a histogram-based feature engineering method, a hybrid physics-informed model, and a fine-tuning strategy, is proposed for online battery degradation diagnosis and knee-onset detection. The hybrid model is first developed and evaluated using a scenario-aware pipeline in protocol cycling scenarios and then fine-tuned to create local models deployed in a dynamic cycling scenario. A 2D histogram-based 17-feature set is found to be the best choice in both source and target scenarios. The fine-tuning strategy is proven to be effective in improving battery degradation mode estimation and degradation phase detection performance in the target scenario. Again, a strong linear correlation was found between the identified knee-onset and knee points. As a result, advanced BMS functions, such as online degradation diagnosis and prognosis, online knee-onset detection and knee prediction, aging-aware battery classification, and second-life repurposing, can be enabled through a battery performance digital twin in the cloud.
@@
2D histogram features to enable online battery diagnosis and knee-onset detection.
@@
A fine-tuning strategy to create local models for online deployment.
@@ 1 Introduction
Contribution: The goal of this work is to fill the gaps indicated above by proposing a transferable physics-informed framework for online battery degradation diagnosis, knee-onset detection, and knee prediction. Specifically, the model takes histograms aggregated from time-series voltage and current data, which are easy to acquire in an onboard BMS. In addition, its performance can be generalized to unseen battery usage scenarios at a satisfactory level using a small amount of labeled data in a target scenario. Our key results and contributions are as follows:
@@
    As a result of curvature-based knee and knee-onset identification, the concept of ”degradation phases” is introduced, with which batteries can be classified according to which degradation phase they are in. Specifically, Phase 1 is defined as the period from the beginning of life to the knee-onset point in which batteries may continue their usage in first-life applications; Phase 2 is defined as the period from the knee-onset point to the knee point in which batteries may either continue their usage in first-life applications or be repurposed to second-life applications in which the knee occurrence may be avoided; Phase 3 is defined as the period beyond the knee point in which batteries may potentially be repurposed to very mild second-life applications in which the knee occurrence can be stopped, or be recycled.
@@
    A transferable physics-informed framework is proposed, which consists of a histogram-based feature engineering method, a hybrid physics-informed model, and a fine-tuning strategy. The 2D histogram-based 17-feature set retains the correlation between current and voltage with its predictive power generalizing across different usage scenarios. The proposed hybrid physics-informed model is the first application of a deep hidden physics model (DeepHPM) for battery degradation mode estimation and an XGBoost model for battery degradation phase detection.
@@
    The hybrid model using a 2D histogram-based 17-feature set is found to be the best choice for estimating battery degradation modes in both source and target scenarios. The fine-tuning strategy was proven to be effective in improving battery degradation mode estimation and phase detection performance in the target scenario. With degradation phases detected with high accuracy, online prediction of battery capacity knee points can also be achieved by leveraging the strong linear correlations identified between knee-onset and knee points.
@@ 2.1 Related definitions
A degradation mode is a degradation mechanism that has a unique and measurable effect on, for example, the capacity, impedance, and open circuit voltage (OCV) of lithium-ion cells, caused by one or multiple interacting degradation mechanisms. Three degradation modes are considered here, loss of lithium inventory (LLI), loss of active material at the negative electrode (LAM_NE), and loss of active material at the positive electrode (LAM_PE). Notably, another degradation mode, i.e., conductivity loss (CL), describes ohmic resistance increase.
Definition 2.3 (Knee [37]).
The capacity knee is defined as the point when ”the capacity slowly declines throughout most of the battery’s life, but begins to decrease rapidly in the latter stages” (see Fig. 1a).
Definition 2.4 (Knee-onset [14]).
The capacity knee-onset is defined as ”the point that marks the beginning of the accelerated degradation rate at which the capacity fade can no longer be approximated as a linear function”. The knee-onset point shows the very first signs of the transition to the accelerated degradation rate, which provides much earlier warning than the knee point, where the accelerated degradation is already taking place (see Fig. 1a).
Definition 2.5 (Degradation phases [38]).
@@ 2.2 Aging-aware battery classification
Lithium-ion batteries are designed, manufactured, and used for a wide range of applications, such as portable electronics, electric vehicles (EVs), and stationary energy storage systems. As a result of complex interactions between multiple degradation mechanisms that are influenced by operating conditions (e.g., charge/discharge current, temperature, state-of-charge (SoC) window, etc.), batteries can exhibit capacity fade curves that are linear, sublinear, superlinear or a combination of the three [21]. In particular, superlinear capacity fade infers a knee on the capacity fade curve, which may significantly shorten a battery’s lifetime and pose safety risks [39]. Therefore, batteries with knee occurrence are recommended to retire immediately for safety concerns [40]. As a result, retired EV batteries with knee occurrence will have no second-life value. However, if we can classify EV batteries during their first-life applications, and then repurpose them to well-controlled second-life applications in which the knee occurrence can be delayed or even avoided, then enormous quantities of EV batteries could potentially have substantial second-life value with guaranteed safety.
@@ 2.3 Histogram-based feature engineering
This histogram-based feature engineering method consists of two steps, i.e., first determining the variable bounds; and then extracting features from time-series usage data. We assume that voltage and current measurements are available for each cell in a battery pack, we will therefore focus on features extracted from the time series of these two variables in this work. Specifically, we first determine the variable bounds, for which a histogram and a cumulative histogram are generated from the measured voltage and current data. An example of the histogram and the cumulative one generated from the voltage data can be seen in Fig. 2a and 2b, respectively. Similarly, an example of the histogram and the cumulative one generated using the current data can be seen in Fig. 3a and 3b, respectively. Each bar in the histograms represents the time spent within a specific voltage or current range. The 1st, 33rd, 67th, and 99th percentiles of each variable are given in Table 1. Based on these bounds, a voltage-based 3-feature set was selected as the optimal set for online capacity estimation by Greenbank and Howey [15]. Although battery capacity fade is attributed to the growth of internal degradation modes, this feature set may not be the best feature set for online battery degradation mode estimation. Moreover, time spent in extreme ranges, for example, ranges outside charge/discharge cut-off voltages, should also be monitored due to its evident degradation effects [43]. As listed in Table 2, five feature sets are generated for online battery degradation mode estimation.
@@
Feature set Input feature 1D voltage-based 3-feature set Time spent between 1st and 33rd voltage percentiles (i.e., $V_{12}$) Time spent between 33rd and 67th voltage percentiles (i.e., $V_{23}$) Calendar time (i.e., $t$) 1D voltage-based 5-feature set Time spent less than 1st voltage percentile (i.e., $V_{01}$) Time spent between 1st and 33rd voltage percentiles (i.e., $V_{12}$) Time spent between 33rd and 67th voltage percentiles (i.e., $V_{23}$) Time spent greater than 67th voltage percentile (i.e., $V_{34}$) Calendar time (i.e., $t$) 1D current-based 3-feature set Time spent between 1st and 33rd current percentiles (i.e., $I_{12}$) Time spent between 33rd and 67th current percentiles (i.e., $I_{23}$) Calendar time (i.e., $t$) 1D current-based 5-feature set Time spent less than 1st current percentile (i.e., $I_{01}$) Time spent between 1st and 33rd current percentiles (i.e., $I_{12}$) Time spent between 33rd and 67th current percentiles (i.e., $I_{23}$) Time spent greater than 67th current percentile (i.e., $I_{34}$) Calendar time (i.e., $t$) 2D current-voltage 17-feature set Time spent less than 1st current percentile and in 4 voltage ranges (i.e., $I_{01}V_{01}$, $I_{01}V_{12}$, $I_{01}V_{23}$, $I_{01}V_{34}$) Time spent between 1st and 33rd current percentiles and in 4 voltage ranges (i.e., $I_{12}V_{01}$, $I_{12}V_{12}$, $I_{12}V_{23}$, $I_{12}V_{34}$) Time spent between 33rd and 67th current percentiles and in 4 voltage ranges (i.e., $I_{23}V_{01}$, $I_{23}V_{12}$, $I_{23}V_{23}$, $I_{23}V_{34}$) Time spent greater than 67th current percentile and in 4 voltage ranges (i.e., $I_{34}V_{01}$, $I_{34}V_{12}$, $I_{34}V_{23}$, $I_{34}V_{34}$) Calendar time (i.e., $t$)
@@ 2.4 Hybrid model architecture design
$u_{i}=f_{i}(t,\bm{x}),\ i=1,2,3,$ (1)
where $t\in\mathbb{R}$ denotes time, $\bm{x}\in\mathbb{R}^{m}$ denotes an input vector and can be one of five feature sets, excluding the calendar time $t$ (see Table 2), and $u_{i}\in\mathbb{R}$ denotes each of the three degradation modes (see Definition 2.2). The degradation mode growth rate is defined as how quickly a degradation mode grows with respect to time, and can be described using a nonlinear partial differential equation (PDE) in the general form
@@
where we have omitted the index $i$ to simplify, $g(\cdot)$ is a nonlinear function of time $t$, the input vector $\bm{x}$, solution $u$, and its derivatives with respect to the input vector, for example, $u_{\bm{x}}=[\partialderivative{u}{x_{1}},\partialderivative{u}{x_{2}},\ldots,\partialderivative{u}{x_{m}}]^{T}$. The function $g(\cdot)$ comprises battery internal degradation dynamics and can represent different forms of degradation, such as linear, sublinear, superlinear, or combination of the three [21]. However, the explicit form of $g(\cdot)$ is unknown and almost impossible to obtain. Inspired by the deep hidden physics model (DeepHPM) proposed by Raissi et al. [44], we approximate the function $f(\cdot)$ in Eqn. (1) and the nonlinear function $g(\cdot)$ in Eqn. (2) with two neural networks (NNs) and define a DeepHPM $\mathcal{H}$ to model battery degradation mode:
@@
where $\mathcal{F}(\cdot)$ denotes the surrogate NN that approximates the hidden solution of the dynamical models, $\mathcal{G}(\cdot)$ denotes the dynamic NN that approximates the battery degradation dynamics, and $\partialderivative{\mathcal{F}(t,\bm{x};\Phi)}{t}$ denotes the partial derivatives of the surrogate NN $\mathcal{F}(\cdot)$ with respect to $t$. Notably, we only consider the first-order partial derivatives in the dynamic NN, to achieve a good trade-off between accuracy and complexity. The parameters $\Phi$ of the surrogate NN $\mathcal{F}(\cdot)$ and $\Theta$ of the dynamic NN $\mathcal{G}(\cdot)$ can be learned by minimizing the sum of squared errors by the loss function
@@
where $n$ is the number of training samples and $\mathcal{H}_{t}=\partialderivative{\mathcal{H}}{t}$. The data loss term $\mathcal{L}_{u}$ aims to find the parameters of the surrogate NN $\mathcal{F}(\cdot)$ so that it fits the data, while the PDE loss term $\mathcal{L}_{\mathcal{H}}$ aims to find the parameters of the dynamic NN $\mathcal{G}(\cdot)$ so that it satisfies the PDE defined by Eqn. (2) at the evaluated points $(t_{i},\bm{x}_{i})$. Moreover, it has been empirically demonstrated that embedding the gradient of the PDE residual into the loss function can further improve the accuracy of the DeepHPM [45]. Therefore, the PDE gradient loss term $\mathcal{L}_{\mathcal{H}_{t}}$ that aims to reduce fluctuations and makes the PDE residual closer to zero is also added here. During the training process, the derivatives of the surrogate NN w.r.t. time $t$ and input $\bm{x}$, and the derivatives of the DeepHPM function w.r.t. time $t$ are evaluated using automatic differentiation [46].
With the availability of battery degradation modes estimated by the DeepHPM, the degradation phase of a battery can be detected. Mathematically, the degradation phase detection can be formulated as a multi-class classification problem to classify each battery into one of the three possible classes (see Subsection 2.2) given its estimated degradation modes and calendar time. As an efficient and scalable machine learning system for tree boosting, proposed by Chen and Guestrin [47], XGBoost has demonstrated outstanding performance in battery state of health estimation [48] [49], and remaining useful life prediction [50]. Thus, it is reasonable to explore its potential for addressing the classification problem here.
@@
$\mathcal{L}^{(t)}=\sum^{n}_{i=1}l(y_{i},\hat{y}^{(t-1)}_{i}+h_{t}(\bm{v}_{i}))+\omega(h_{t})+\mathrm{constant},$ (8)
where $n$ is the number of training samples, $y_{i}$ is the value of the $i$-th sample, $\hat{y}^{t-1}_{i}$ is the prediction of the $i$-th sample up to the (t1)(t-1)-th iteration, $h_{t}(\bm{v}_{i})$ is the output of the $t$-th tree for the $i$-th sample. In our case, the XGBoost model is to approximate a function that predicts the class $y\in\{1,2,3\}$ given an input vector $\bm{v}=[u_{1},u_{2},u_{3},t]^{T}$. Note that $u_{1}$, $u_{2}$, and $u_{3}$ represent three ”true” degradation modes governed by three Eqn. (1), respectively. Thus, the general loss function in Eqn. (8) must be modified to account for multi-class classification as
$\mathcal{L}^{(t)}=\sum^{n}_{i=1}\sum^{3}_{k=1}l(y^{k}_{i},\hat{y}^{k,(t-1)}_{i}+h^{k}_{t}(\bm{v}_{i}))+\omega(h_{t})+\mathrm{constant},$ (9)
@@
The resulting hybrid physics-informed model is illustrated in Fig. 8A.